Question 8 (2.5 points) A hedge fund leverages its $100 million of investor capital by a factor of four (borrows $300 million) and invests it into a portfolio of junk bonds yielding 12%. If its borrowing costs are 9%, what is the yield on investor capital? 1) 14% 2) 18% 3) 26% 4) 21%
Answer #1

interest to be paid = 300*9% = 27 million

return received = 12% of 400 = 48 million

return on investor capital = (47 - 27)/100 = 21%